Find the real solutions of the equation by graphing.
x2 + 2x + 2 = 0

Found 2 solutions by ewatrrr, jim_thompson5910:
Answer by ewatrrr(24785)

 
Hi,
y = x^2 + 2x + 2 Pt(0,2) on the curve
y = x^2 + 2x +1 -1 + 2
y = (x + 1)^2 + 1 parabola, V(-1,1) line of symmetry x = -1
Graph, NO real solutions


Answer by jim_thompson5910(35256)

Use a graphing calculator to get



The green graph NEVER crosses the x axis. So there are NO real solutions.